OpenStudy (anonymous):

Multiply and then simplify the radical. 8√ times √3

OpenStudy (yamyam70):

\[8\sqrt{?} \times \sqrt{3} is this \it ? \]

OpenStudy (yamyam70):

is that it?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

oh, no wait

OpenStudy (debbieg):

There is nothing under that first radical sign?? lol

OpenStudy (anonymous):

its √8 times √3

OpenStudy (yamyam70):

we just multiply the variables, and we take the square root of the product :)

OpenStudy (anonymous):

answers are: a) 2√6 b) √24 c) 6√2 d) 4√6

OpenStudy (anonymous):

how do u do that O_o

OpenStudy (yamyam70):

/8 times /3 = /24 , get it?

OpenStudy (yamyam70):

then you factor 24 with a perfect square :)

OpenStudy (anonymous):

O.O thats all you have to do!??!

OpenStudy (yamyam70):

yup :))

OpenStudy (anonymous):

LOL

OpenStudy (yamyam70):

do you get it the whole process ? :)

OpenStudy (anonymous):

i guess so, thanks =)